If a unique solution does not exist,how will I know if the system is inconsistent or dependent? How do I solve the following system?
5x+2y=30
x-4y=17

Answer by stanbon(48558)

5x+2y=30
x-4y=17
----------
The system is "dependent" if one equation is a multiple of the other.
The system is "inconsistent" if when you assume there is a solution you
run into a contrdiction like 3=0.
---------------
To "solve" your system by substitution, solve the 2nd equation for "x":
x=4y+17
Substitute that into the 1st equation:
5(4y+17)+2y=30
20y+85+2y=30
22y=-55
y=-5/2
---------
Substitute that into x=4y+17 to solve for x:
x=4(-5/2)+17
x=7
-----------
We assumed there was a solution and found out there is.
The solution is (7,-5/2)
Therefore the system is said to be consistent.
---------
Cheers,
Stan H.